A 10 by 10 array of horizontal tubes of 1.27 cm diameter is exposed to pure steam at atmospheric



pressure. If the tube wall temperature is 98°C, estimate the mass of steam condensed assuming a



tube length of 1.5 m

1
Expert's answer
2022-01-31T15:06:01-0500

10*10*1.27=127cm3

"m=v*T*l"

"=127*98*1.5"

=18669g

=18.669kg